5km in 20 min as a rate


1
Expert's answer
2020-12-15T01:25:59-0500

Distance = 5 km


1km = 1000metres

5 km = ?

5(1000) = 5000metres


Time = 20 minutes


1minute = 60seconds

20minutes = ?

20(60) =1200secomds


Speed = D/T


Therefore = ( 5000/1200) m/s


= 50/12 m/s= 4.2 m/s